The mean score on the the exam was 63%. The Standard deviation was 12%. Find the smallest interval about the mean that contains at least 82% of the scores.
Answer by ewatrrr(24785)

 
Hi,
μ = 63% and σ = 12%
Find the smallest interval about the mean that contains at least 82% of the scores
(100-82)/2 = 18/2 = 9% on either side
z = 1.341
1.341 = (x-63)/12
1.341*12 + 63 = 79
79-63 = 13
82% of the scores will be between 50%-79%